OpenStudy (anonymous):

Find the exact value of arccos (sqrt(2)/(2))

OpenStudy (anonymous):

arctan( sqrt(3)/3 ) = arctan( 3 / sqrt(3) ) because ( sqrt(3)/3 )*( sqrt(3) )/sqrt(3) ) = 3/( 3*sqrt(3) ) = 1/sqrt(3) Use the 1 - sqrt(3) - 2 triangle to figure this out. arctan( 1/sqrt(3) ) = pi/6 = 30 degrees pi/6 = 30 degrees because pi/6 * 180 degrees/pi = 30 degrees

OpenStudy (anonymous):

I dont realy know what is it about

OpenStudy (anonymous):

The answer is 45 degrees or pi/4. arccos is a function which asks you which angle x has the cos(x)=some value. In this case, they're just asking you which angle has the cos value of sqrt(2)/2.